Suppose New Zealand is open to free trade in the world market for wheat. Because of New Zealands small size, the demand forNow suppose the New Zealand government decides to impose a tariff of $60 on each imported ton of wheat. After the tariff, thePRICE (Dollars per ton) Government Revenue 0 20 40 160 180 200 60 80 100 120 140 QUANTITY (Tons of wheat) DWL Complete the fofill in the blank

1)increase/decrease

2)increase/decrease

3)gain/loss
Answer #1

Domestic Demand Domestic Supply 8 PRICE (Dollars per ton) 0 20 40 160 180 200 60 80 100 120 140 QUANTITY (Tons of wheat)

Import = 160-40 =120 tons

With tariff, the price increases to 310 and Imports = 120-80 = 40 tons

Domestic Demand Domestic Supply World Price Plus Tariff PRICE (Dollars per ton) Government Revenue 190 0 20 40 160 180 200 DW

Free trade Tariff
CS 0.5*160*(490-250) = 19200 0.5*120*(490-310) = 10800
PS 0.5*40*(250-190) = 1200 0.5*80*(310-190) = 4800
GR 0 40*60 = 2400

CS decreases by 19200-10800 = 8400

PS increases by 4800-1200 = 3600

GR = 2400

Net welfare effect is a loss of 2400